Let T: R^2 to R^3 be defined by (a, b)=(a+b, a-2b, 3a+b). Show that T is nonsingular. Hence find T inverse.
1
Expert's answer
2020-12-17T03:27:07-0500

"T(a,b)=(a+b,a-2b,3a+b)"


Let "T(a,b)=0\\" then (a+b,a-2b,3a+b)=0


"\Rightarrow a=0, b=0"

Therefore "(a,b)=(0,0)."

Hence T is nonsingular.


Therefore T"^{-1}" exists from "R^{3}" to "R^{2}".


Now let, "(a+b,a-2b,3a+b)=(x,y,z)"


"\Rightarrow" a+b=x

a-2b=y

3a+b=z


Now equating these equations one can get "a=\dfrac{2z+y}{7}, b=\dfrac{z-3y}{7}"


Therefore the inverse of T is "T^{-1}: R^{3}\rightarrow R^{2}\\
T^{-1}(x,y,z)=(\dfrac{2z+y}{7}, \dfrac{z-3y}{7})"